# Shuffle Mamba: State Space Models with Random Shuffle for Multi-Modal Image Fusion

### v1

@@ Abstract
Multi-modal image fusion integrates complementary information from different modalities to produce enhanced and informative images. Although State-Space Models, such as Mamba, are proficient in long-range modeling with linear complexity, most Mamba-based approaches use fixed scanning strategies, which can introduce biased prior information. To mitigate this issue, we propose a novel Bayesian-inspired scanning strategy called Random Shuffle, supplemented by an theoretically-feasible inverse shuffle to maintain information coordination invariance, aiming to eliminate biases associated with fixed sequence scanning. Based on this transformation pair, we customized the Shuffle Mamba Framework, penetrating modality-aware information representation and cross-modality information interaction across spatial and channel axes to ensure robust interaction and an unbiased global receptive field for multi-modal image fusion. Furthermore, we develop a testing methodology based on Monte-Carlo averaging to ensure the model’s output aligns more closely with expected results. Extensive experiments across multiple multi-modal image fusion tasks demonstrate the effectiveness of our proposed method, yielding excellent fusion quality over state-of-the-art alternatives. Code will be available upon acceptance.
Introduction
@@
Multi-modal image fusion, a fundamental task in computer vision, involves extracting and integrating valuable information from images of the same scene captured by different imaging modalities. This task aims to create a single composite image with a more comprehensive and informative representation, with typical applications including pan-sharpening and medical image fusion (MIF). In the context of pan-sharpening, satellites are limited by sensors, which can only capture low-resolution multi-spectral (LRMS) and panchromatic (PAN) images. Specifically, PAN images offer superior spatial details but limited spectral resolution, while MS images provide abundant spectral resolution but lack spatial clarity. Through integrating the complementary information from both MS and PAN images into a composite representation, we can achieve an effective balance between spatial and spectral resolution. Analogously, in the realm of MIF, various imaging technologies capture distinct types of information. For instance, Computed Tomography (CT) images deliver detailed insights into bones and high-density tissues, while Magnetic Resonance Imaging (MRI) offers higher-resolution images with rich soft tissue details. In virtue of this complementary information from various modalities, MIF can overcome the limitations of single-modality images, resulting in a more comprehensive and detailed representation for modern medical diagnosis.
Figure 1: Visualization of Effective Receptive Fields (ERFs) for Conv, Self-Attention and our Method. A larger ERF is indicated by a more extensively distributed dark area.
@@
In recent years, the prosperous advancement of deep neural networks (DNNs) has led to the development of numerous DNN-based multi-modal image fusion methods. In pan-sharpening, the pioneering work PNN (Masi et al. 2016) employed a simple three-layer neural network to achieve remarkable results which was previously deemed unattainable, highlighting the superior learning capabilities of deep learning. After that,increasingly complex and deeper architectures have emerged (Zhou et al. 2022b; Yang et al. 2023), delivering excellent visual performance. Despite these advancements, existing multi-modal fusion methods face common limitations. Convolutional neural network (CNN)-based approaches often struggle to establish global receptive fields. While transformers address this issue through self-attention mechanisms, they introduce significant challenges related to quadratic computational complexity. Nowadays, structured state-space models have gained considerable attention for their computational efficiency and principled ability to model long-range dependencies. However, their selective scanning mechanism can introduce biased priors when processing 2D images. To overcome these challenges, it is reasonable for us to design a novel sequential scanning method with its application framework.
@@
Our Motivation. Global modeling capabilities are essential in image restoration tasks because one of the core aspects of image restoration is to find useful information within the image to compensate for the missing data in the current patch. For a long time, CNN and ViT have been dominant architectures in computer vision, each with distinct advantages and limitations. CNNs are constrained by local receptive fields, which hinder their ability to model long-range dependencies. In contrast, ViT uses the self-attention mechanism to access a global receptive field but are burdened by quadratic computational complexity. Recently, structured state-space models have demonstrated enhanced capabilities in capturing long-term dependencies in sequence data while maintaining linear time complexity. Notably, Mamba has achieved significant strides in reducing inference latency and improving overall performance through selective state spaces and hardware-aware algorithms. With the introduction of vmamba (Liu et al. 2024) and Vision Mamba (Zhu et al. 2024), there has been a growing interest in applying state-space models to visual tasks.
However, most current Mamba-based methods employ unidirectional SSMs, which endure certain limitations in their receptive field range due to their scanning approaches. Specifically, the receptive field is large for the initial portion of input patches but becomes significantly smaller for the latter portion, potentially compromising the model’s global modeling capabilities. Moreover, dissimilar to the dependencies between token orders in language modeling, the non-causal nature of 2-D spatial information in images presents a considerable challenge for simple sequential scanning methods. Traditional strategies, such as flattening image patches and scanning them sequentially, may introduce biased local 2D dependencies, thereby undermining the model’s ability to effectively identify spatial relationships.
To address these challenges, we propose a new sequence scanning method called Random Shuffle Scanning. Figure 1 shows the visualization of the effective receptive field (ERF) of three methods. Compared to the Conv method, the Self-Attention method’s advantage lies in achieving a global receptive field. However, our method provides an overall unbiased global receptive field, enabling the network to discard fixed local priors and focus more effectively on what needs to be learned. For the sequential input image patches, we first apply position encoding, followed by a random shuffle of the patches before they are processed by the Mamba block for long-range dependency modeling. The random shuffle approach eliminates the preconception between local and global 2D dependencies in mathematical expectation, enabling the model to access an unbiased prior and ultimately establish a more consistent and global receptive field. Recognizing that shuffling the spatial position of the image patches may disrupt semantic information, we implement a corresponding inverse transformation to accurately restore the sequence order of the input patches after passing through the Mamba block. This information-preserving transformation pair underpins the key components of our model: Random Mamba Block, Random Channel Interactive Mamba Block, and Random Modal Interactive Mamba Block. Inspired by Dropout, we use Monte Carlo averaging to approximate the expected output, ensuring that the actual output at test time closely matches the anticipated results.
@@
Our contributions can be summarized as follows: 1) We design the Shuffle Mamba framework, where the random shuffle operation in the key component provides an expected unbiased global receptive field without increasing any parameters. 2) We develop a specific strategy for training and testing this framework. During training, each input is independently scanned using a random shuffle operation. During testing, we use Monte Carlo averaging to estimate the output of each Mamba block. 3) Extensive experiments on two prominent multi-modal image fusion tasks demonstrate that our method accomplishes excellent performance in both quantitative assessments and visual quality.
Related Works
State Space Model
The State Space Model (SSM), originating from control theory, has found extensive applications in deep learning due to its remarkable ability to model long-range dependencies. Initially, the S4 (Gu, Goel, and Ré 2021) model introduced the concept of the SSM, effectively reducing the computational and memory requirements associated with state representation while enabling global information modeling. Building upon S4, the S5 (Smith, Warrington, and Linderman 2022) model features MIMO structure and efficient parallel scanning strategy to enhance performance without significantly increasing computational demands. Furthermore, the H3 (Mehta et al. 2022) model further refines these methods, achieving competitive performance and efficiency comparable to the transformer in language modeling tasks.
@@
Recently, Mamba (Gu and Dao 2023) has significantly improved inference speed and performance metrics through selective state spaces and hardware-aware algorithms. The introduction of Vmamba (Liu et al. 2024) and Vision Mamba (Zhu et al. 2024) has brought attention to the application of SSM in vision tasks. However, most existing SSM-based vision models (Zhu et al. 2024; Ma, Li, and Wang 2024; Zheng and Wu 2024) employ a fixed scanning strategy, which may introduce preconceptions, particularly in low-level vision tasks (Xiao et al. 2023). Specifically, this fixed order of choosing image patches can cause the model to gradually dismiss previous input sequences while processing the current patch, thereby compromising its capability to model global information. To alleviate this challenge, Vmamba (Liu et al. 2024) introduced the CSM, which scans image pixels from various directions such as top-left, bottom-right, top-right, and bottom-left. Building on this idea, RSM (Zhao et al. 2024) devised the OSSM to flatten image patches into sequences in eight directions, enhancing the network’s capability to capture and model large-scale spatial features. Additionally, LocalMamba (Huang et al. 2024) applies the windowed selective scan to ensure a harmonious integration of global and local visual cues. RSMamba (Chen et al. 2024) incorporates dynamic multi-path activation mechanisms to model non-causal data. However, these methods do not simultaneously consider the integrity of the image’s structure and the randomness of the pixels during scanning processing, and are essentially still special fixed strategies. To address this limitation, we introduced the Shuffle Mamba framework, which includes random shuffle and inverse operations to obtain a global receptive field without bias in expectations.
Method
Preliminaries
@@
Network Architecture
The proposed Shuffle Mamba framework consists of three functional components: the Random Mamba block (RM block), the Random Channel Interactive Mamba block (RCIM block), and the Random Modal Interactive Mamba block (RMIM block). The overall workflow is illustrated in Figure 2. Assuming that the input images with different modalities are ${\mathbf{M}_{up}}$ and ${\mathbf{P}_{in}}$, where ${\mathbf{M}_{up}}$ serve as the upsampled ${\mathbf{M}_{in}}$, we first use convolutional layers to project the images into the feature space. Given the limited receptive field of the convolutional layers, which makes capturing global features challenging, we perform patch embedding and send the resulting patches to the RM block for global feature extraction. This process yields the global modality-specific features $\mathbf{F}_{n}^{m}$ and $\mathbf{F}_{n}^{p}$, where $PE=PatchEmbed$:
$\mathbf{F}_{0}^{m},\mathbf{F}_{0}^{p}=PE(\phi({\mathbf{M}_{up}})),PE(\phi({\mathbf{P}_{in}}))$ (8)
$\mathbf{F}_{n}^{m}={{\psi}_{1n}}\cdot\cdot\cdot({{\psi}_{11}}({{\psi}_{10}}(\mathbf{F}_{0}^{m}))$ (9)
$\mathbf{F}_{n}^{p}={{\psi}_{2n}}\cdot\cdot\cdot({{\psi}_{21}}({{\psi}_{20}}(\mathbf{F}_{0}^{p}))$ (10)
$\mathbf{F}_{k+j}^{m}={{\theta}_{j}}\cdot\cdot\cdot({{\theta}_{1}}({{\theta}_{0}}(\mathbf{F}_{k}^{m},\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p})$ (11)
${\mathbf{H}_{out}}=\phi(reshape(\mathbf{F}_{k+j}^{m}))+{\mathbf{M}_{up}}$ (12)
@@
Key Components
Random Shuffle Scanning.
Mamba was originally designed to adapt to the modeling of language sequences. We devised a Random Shuffle Scanning approach to access unbiased local and global dependencies in 2D image while ensuring a global receptive field similar to self-attention. As shown in Figure 3, for sequentially input 2D image patches, we first apply depth-wise convolution for position modeling. The image patches are then randomly shuffled and sent to the Mamba block for long-range dependency modeling. This strategy allows the Mamba block to simulate interactions between adjacent patches with equal probability, enabling the network to learn and model from an unbiased prior effectively. Additionally, since the relative positions of the patches are crucial for reconstructing semantic information, the output image must be accurately aligned with the input based on the inverse shuffle. For this reason, the random shuffle and its inverse operation constitute an information-lossless transformation pair.
Random Mamba Block.

### v2

# arXiv:2409.01728v2[cs.CV] 27 Feb 2026

@@ Abstract
Multi-modal image fusion integrates complementary information from different modalities to produce enhanced and informative images. Although State-Space Models, such as Mamba, are proficient in long-range modeling with linear complexity, most Mamba-based approaches use fixed scanning strategies, which can introduce biased prior information. To mitigate this issue, we propose a novel Bayesian-inspired scanning strategy called Random Shuffle, supplemented by a theoretically feasible inverse shuffle to maintain information coordination invariance, aiming to eliminate biases associated with fixed sequence scanning. Based on this transformation pair, we customized the Shuffle Mamba Framework, penetrating modality-aware information representation and cross-modality information interaction across spatial and channel axes to ensure robust interaction and an unbiased global receptive field for multi-modal image fusion. Furthermore, we develop a testing methodology based on Monte-Carlo averaging to ensure the model’s output aligns more closely with expected results. Extensive experiments across multiple multi-modal image fusion tasks demonstrate the effectiveness of our proposed method, yielding excellent fusion quality compared to state-of-the-art alternatives. The code is available at https://github.com/caoke-963/Shuffle-Mamba.
Index Terms:
State-space model, Multi-modal Image fusion, Pan-sharpening
I Introduction
@@
Multi-modal image fusion, a fundamental task in computer vision, involves extracting and integrating valuable information from images of the same scene captured by different imaging modalities [29, 27]. This task aims to create a single composite image with a more comprehensive and informative representation, with typical applications including pan-sharpening and medical image fusion (MIF). In the context of pan-sharpening, satellites are limited by sensors, which can only capture low-resolution multi-spectral (LRMS) and panchromatic (PAN) images. Specifically, PAN images offer superior spatial details but limited spectral resolution, while MS images provide abundant spectral resolution but lack spatial clarity. By integrating the complementary information from both MS and PAN images into a composite representation, we can achieve an effective balance between spatial and spectral resolution. Analogously, in the realm of MIF, various imaging technologies capture distinct types of information. For instance, Computed Tomography (CT) images deliver detailed insights into bones and high-density tissues, while Magnetic Resonance Imaging (MRI) offers higher-resolution images with rich soft tissue details. In virtue of this complementary information from various modalities, MIF can overcome the limitations of single-modality images, resulting in a more comprehensive and detailed representation of modern medical diagnosis.
Figure 1: Visualization of Effective Receptive Fields (ERFs) and ERF Intensity Distributions for Conv, Self-Attention, Pan-Mamba, and Our Method. The upper part displays ERF visualizations, where larger, darker regions indicate broader receptive fields. The lower part presents the histogram of ERF intensities, with the x-axis representing ERF strength (lower values correspond to darker colors and stronger ERF responses) and the y-axis denoting pixel probability density. Additionally, the mean and standard deviation of ERF intensities are reported to provide a comprehensive comparison.
@@
In recent years, the prosperous advancement of deep neural networks (DNNs) has led to the development of numerous DNN-based multi-modal image fusion methods. In pan-sharpening, the pioneering work PNN [38] employed a simple three-layer neural network to achieve remarkable results that were previously deemed unattainable, highlighting the superior learning capabilities of deep learning. After that, increasingly complex and deeper architectures emerged [75, 66, 69], delivering excellent visual performance. Despite these advancements, existing multi-modal fusion methods face common limitations. Beyond improving fusion quality in standard scenarios, a growing trend in the field is to address image fusion in complex and adverse conditions. Recent studies have proposed unified frameworks to handle multiple complex scenes [24] and developed large-scale benchmarks to facilitate all-weather multi-modality fusion [26]. Furthermore, incorporating global and local text perception has proven effective for fusion in adverse weather [25]. These works highlight the increasing demand for robust perception capabilities in real-world applications. Convolutional neural network (CNN)-based approaches often struggle to establish global receptive fields. While transformers address this issue through self-attention mechanisms, they introduce significant challenges related to quadratic computational complexity. Nowadays, structured state-space models have gained considerable attention for their computational efficiency and principled ability to model long-range dependencies. However, their selective scanning mechanism can introduce biased priors when processing 2D images. To overcome these challenges, it is reasonable for us to design a novel sequential scanning method with its application framework.
@@
Our Motivation. Global modeling capabilities are essential in image restoration tasks because one of the core aspects of image restoration is to find useful information within the image to compensate for the missing data in the current patch. For a long time, CNN and ViT have been dominant architectures in computer vision, each with distinct advantages and limitations. CNNs are constrained by local receptive fields, which hinder their ability to model long-range dependencies. In contrast, ViT uses the self-attention mechanism to access a global receptive field but is burdened by quadratic computational complexity. Structured state-space models have recently demonstrated enhanced capabilities in capturing long-term dependencies in sequence data while maintaining linear-time complexity. In particular, Mamba has made significant strides in reducing inference latency and improving overall performance through selective state spaces and hardware-aware algorithms. With the introduction of VMamba [33] and Vision Mamba [76], there has been a growing interest in applying state-space models to visual tasks.
Most existing Mamba-based vision models adopt unidirectional or fixed scanning strategies to convert 2D spatial data into 1D sequences, which introduces fundamental limitations in receptive field modeling. Due to the sequential nature of state space models, earlier tokens typically benefit from broader receptive fields, while later tokens suffer from reduced context, leading to unbalanced global dependency modeling. Moreover, unlike the inherently causal structure of language, visual data is spatially non-causal and two-dimensional. Flattening image patches and scanning them sequentially disrupts spatial continuity, introduces orientation-specific biases, and hampers the modeling of local context. Prior studies, such as DgMamba [35] and EAMamb [30], have shown that fixed scanning breaks adjacent pixel relations, while MambaIR [10] and MS-VMamba [46] demonstrate that such strategies lead to uneven receptive field distributions and long-range information decay. These issues persist even with multi-directional scanning, indicating that deterministic scanning orders inherently limit the spatial modeling capacity of Mamba-based frameworks.
To address these challenges, we propose a new sequence scanning method, termed Random Shuffle Scanning. Figure 1 visualizes the effective receptive fields (ERFs) of four representative approaches. Compared with convolutional models, self-attention achieves global perception but at the cost of quadratic computational complexity. Pan-Mamba achieves global modeling with linear complexity; however, its deterministic scanning order introduces structural bias, leading to an overemphasis on specific image patterns, such as horizontal stripes. In contrast, the proposed Shuffle Mamba establishes a balanced global receptive field with linear-time complexity, effectively removing fixed local priors and allowing the model to focus on task-relevant regions adaptively. To further investigate the receptive field distribution, the lower part of Figure 1 presents histograms of pixel-wise ERF intensities. The horizontal axis denotes gradient-based response intensity, while the vertical axis represents probability density. Compared with other methods, Shuffle Mamba exhibits a left-shifted distribution with a smaller mean value, corresponding to darker regions in the visualization and stronger activation responses. These results indicate that our randomized scanning promotes stronger global correlations and more uniform dependency modeling across spatial locations, effectively mitigating local bias and achieving theoretically unbiased global modeling. For the sequentially input image patches, we first apply position encoding and then randomly shuffle the patches before processing them through the Mamba block for long-range dependency modeling. This random shuffle operation removes the deterministic correlation between local and global 2D dependencies in expectation, enabling the model to access an unbiased prior and achieve consistent global perception. Recognizing that spatial shuffling may disrupt semantic consistency, a corresponding inverse transformation is applied to restore the original order of input patches after Mamba processing. This information-preserving shuffle–inverse pair serves as the foundation for the three key modules of our framework: the Random Mamba Block, Random Channel Interactive Mamba Block, and Random Modal Interactive Mamba Block. Furthermore, inspired by Dropout, we employ Monte Carlo averaging during inference to approximate the expected output, ensuring that the final prediction aligns closely with the theoretical expectation.
Our contributions can be summarized as follows:
  • We design the Shuffle Mamba framework, where the random shuffle operation in the key component provides an expected unbiased global receptive field without increasing any parameters.
@@
  • We develop a specific strategy for training and testing this framework. During training, each input is independently scanned using a random shuffle operation. During testing, we use Monte Carlo averaging to estimate the output of each Mamba block.
  • Extensive experiments on two prominent multi-modal image fusion tasks demonstrate that our method accomplishes excellent performance in both quantitative assessments and visual quality.
II Related Work
II-A Pan-sharpening
Traditional pan-sharpening methods can be categorized into three main groups: Component Substitution (CS)-based methods [19, 28, 7], Multi-Resolution Analysis (MRA)-based methods [44, 41], and Variational Optimization (VO)-based methods [5, 42]. CS-based approaches extract the spatial details of the PAN image in the transform domain and inject them into the spatial components of the LRMS image. Conversely, MRA-based methods rely on a multi-resolution decomposition to derive the spatial information from the MS image and replace it with the content of the PAN image. VO-based models incorporate prior knowledge to formulate the fusion process and solve it iteratively. However, despite their diverse strategies, these traditional methods often experience significant spectral or spatial distortions due to inappropriate transformation schemes and ill-posed solution conditions.
Inspired by SRCNN [4] in image super-resolution, PNN [38] adopted a simple model consisting of a three-layer convolutional neural network and achieved remarkable results, which has been seen as pioneering work for deep learning in pan-sharpening. This paved the way for many subsequent DL-based approaches with more complex structures. For instance, PanNet [67] conducts high-frequency filtering and residual connections to preserve detailed information and spectrum in the image. MSDCNN [68] incorporates a multi-scale and multi-depth network to extract various-level features. Meanwhile, SFINet [75] and MSDDN [14] leverage the Fourier transform and specially designed dual-domain structures to simultaneously capture spatial and frequency domain features, enabling deeper learning of high-frequency information. Transformer architecture has also been introduced to pan-sharpening [73], which employs self-attention mechanisms to enhance fusion quality. Additionally, considering the ill-posed property of the pan-sharpening problem, PanFlowNet [66] adopts a flow-based model to learn the conditional distribution of the fused image, which can generate multiple variations of HRMS images during inference. More recently, recognizing the limitations of previous methods in global information modeling and cross-modal information fusion, Pan-mamba [12] customizes the core components based on the state space model, yielding excellent results in feature extraction and fusion.
II-B Medical Image Fusion
Medical image fusion (MIF) plays a crucial role in improving disease diagnosis by combining essential information from different image sources into a single fused image. Since single-modality images capture limited information, MIF has become a vital tool to integrate diverse image features, facilitating more accurate and efficient diagnoses. Early methods for MIF predominantly relied on techniques such as spatial domain processing [45], transform domain approaches [22], and sparse representations [50]. However, these traditional methods often struggled to capture deep semantic information and required manually designed fusion rules tailored to specific scenarios, which limited their generalizability. In recent years, deep learning techniques have gained increasing attention in the field of MIF. Liu et al. [32] were among the first to apply convolutional neural networks (CNNs) to MIF tasks, directly learning the mapping from source images to fused outputs. Building on this, Zero [20] utilized pre-trained CNNs to extract deep feature maps and adjust fusion weights by comparing these maps. To tackle the challenge of modeling global context in image fusion, MSRPAN [6] introduced a multi-scale pyramid network, residual connections, and attention mechanisms. This approach achieved superior visual quality and quantitative results by controlling the energy ratio during feature fusion. Furthermore, CDDFuse [71] introduced a correlation-driven Transformer-CNN dual-branch feature decomposition network, effectively separating modality-specific and modality-shared features. Inspired by the mixture of experts (MoE) model, TC-MoA [77] utilized task-specific routing strategies and customized adapters, achieving excellent performance across various image fusion tasks within a unified framework.
@@ II-B Medical Image Fusion
II-C State Space Model
The State Space Model (SSM), originating from control theory, has found extensive applications in deep learning due to its remarkable ability to model long-range dependencies. Initially, the S4 [9] model introduced the concept of the SSM, effectively reducing the computational and memory requirements associated with state representation while enabling global information modeling. Building upon S4, the S5 [47] model features a MIMO structure and an efficient parallel scanning strategy to enhance performance without significantly increasing computational demands. Furthermore, the H3 [39] model further refines these methods, achieving competitive performance and efficiency comparable to the transformer in language modeling tasks.
Recently, Mamba [8] has significantly improved inference speed and performance metrics through selective state spaces and hardware-aware algorithms. The introduction of Vmamba [33] and Vision Mamba [76] has brought attention to the application of SSM in vision tasks. However, most existing SSM-based vision models [76, 37, 72] employ a fixed scanning strategy, which may introduce preconceptions, particularly in low-level vision tasks [57]. Specifically, this fixed order of selecting image patches can cause the model to gradually disregard previous input sequences while processing the current patch, thereby compromising its ability to model global information. To alleviate this challenge, Vmamba [33] introduced the CSM, which scans image pixels from various directions such as top-left, bottom-right, top-right, and bottom-left. Building on this idea, RSM [70] introduced the OSSM to transform image patches into sequences in eight directions, thereby enhancing the network’s ability to capture and model large-scale spatial features. Additionally, LocalMamba [16] applies the windowed selective scan to ensure a harmonious integration of global and local visual cues. RSMamba [3] incorporates dynamic multi-path activation mechanisms to model non-causal data. However, these methods do not simultaneously consider the integrity of the image’s structure and the randomness of the pixels during scanning processing, and are essentially still special fixed strategies.
To address this limitation, we introduced the Shuffle Mamba framework, which employs random shuffling and inverse operations to achieve a global receptive field without introducing expected biases. As a result, our approach not only advances pansharpening and medical image fusion but also demonstrates significant potential for a wider range of low-level vision tasks, including infrared and visible image fusion [2, 1, 17], pan-denoising [63, 64], and multi-modal UAV image fusion [18].
III Methods
In this section, we introduce the foundational concepts of state-space models. Next, we describe the training strategy and module design underlying the proposed random shuffle framework. Finally, we detail the Monte Carlo averaging approach used during testing and discuss the design of the loss function tailored to different tasks.
III-A Preliminaries
@@
III-B Network Architecture
The proposed Shuffle Mamba framework consists of three functional components: the Random Mamba block (RM block), the Random Channel Interactive Mamba block (RCIM block), and the Random Modal Interactive Mamba block (RMIM block). The overall workflow is illustrated in Figure 2. Assuming that the input images with different modalities are $\mathbf{M}_{up}$} and $\mathbf{P}_{in}$}, where $\mathbf{M}_{up}$} serve as the upsampled $\mathbf{M}_{in}$}, we first use convolutional layers to project the images into the feature space. Given the limited receptive field of the convolutional layers, which makes capturing global features challenging, we perform patch embedding and send the resulting patches to the RM block for global feature extraction. This process yields the global modality-specific features $\mathbf{F}_{n}^{m}$ and $\mathbf{F}_{n}^{p}$, where $PE$ means the process of $PatchEmbed$, $\phi$ denotes the initial convolution operation, $\psi$ corresponds to the Random Mamba module, and the subscripts indicate two branches and their sequential modules.
$\mathbf{F}_{0}^{m},\mathbf{F}_{0}^{p}=PE(\phi({\mathbf{M}_{up}})),PE(\phi({\mathbf{P}_{in}})),$ (8)
$\mathbf{F}_{n}^{m}={{\psi}_{1n}}\cdot\cdot\cdot({{\psi}_{11}}({{\psi}_{10}}(\mathbf{F}_{0}^{m})),$ (9)
$\mathbf{F}_{n}^{p}={{\psi}_{2n}}\cdot\cdot\cdot({{\psi}_{21}}({{\psi}_{20}}(\mathbf{F}_{0}^{p})).$ (10)
$\mathbf{F}_{k+j}^{m}={{\theta}_{j}}\cdot\cdot\cdot({{\theta}_{1}}({{\theta}_{0}}(\mathbf{F}_{k}^{m},\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),$ (11)
${\mathbf{H}_{out}}=\phi(reshape(\mathbf{F}_{k+j}^{m}))+{\mathbf{M}_{up}}.$ (12)
III-C Key Components
III-C1 Random Shuffle Scanning
Mamba was originally designed to adapt to the modeling of language sequences. We devised a Random Shuffle Scanning approach to access unbiased local and global dependencies in the 2D image while ensuring a global receptive field similar to self-attention. As shown in Figure 3, for sequentially input 2D image patches, we first apply depth-wise convolution for position modeling. The image patches are then randomly shuffled and sent to the Mamba block for long-range dependency modeling. This strategy allows the Mamba block to simulate interactions between adjacent patches with equal probability, enabling the network to learn and model from an unbiased prior effectively. Additionally, since the relative positions of the patches are crucial for reconstructing semantic information, the output image must be accurately aligned with the input based on the inverse shuffle. For this reason, the random shuffle and its inverse operation constitute an information-lossless transformation pair.
III-C2 Random Mamba Block